From: STOSC::SMM "On-Duty SMM x5074" 8-JAN-1993 11:35:12.27 To: @[.DIS]HST_STAT CC: Subj: HST Status Report, 01/07 - 01/08. HUBBLE SPACE TELESCOPE DAILY REPORT #799 PERIOD COVERED: 0800 EST 01/07/93 - 0800 EST 01/08/93 Daily Status Report as of 008/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/@ 3825 (Novae in Outburst in the Magellanic Clouds) The High Resolution Spectrograph was used to make 14 target of opportunity spectral observations of Magellanic Cloud Novae which is currently in its nebular phase. The observations were received, and where possible displayed. The third through the eighth observations did not occur because of a failed baseline guide star acquisition. The wave length calibrations and the deflection calibrations were reported to be okay. The spectral observations were weak. 1.2 completed FOS/BL 4495 (GIMP Test) A special test was conducted to test the Geomagnetic Induced Movement Problem (GIMP) correction mechanism in the Faint Object Spectrograph. The first observations was received and displayed. The first GIMP test software was gone this week. The results of the test on the blue side were encouraging, the expected X and Y deflections were observed in the telemetry. The STScI analysis of the data showed a marked improvement in the blue side's resolution. The lines showed the typical S-curve as typical of the GIMP. The next two observations generated many status buffer error messages, and an HSTAR was generated. The next nine observations showed a significant amount of drift during the spacecraft's orbital motion. 1.3 Uplinked Updated Ephemris Tables, Per MOSES 2871 1.4 Completed Two Sets of FOC/96 3801 (Search for QSOs suitable for Observation) The Faint Object Camera (f/96) was used to make observations of quasi-stellar object 0830+1009 to measure the level of its ultraviolet flux. The observations were reported to be blank, with no target visible in the image. 1.5 Completed HSP/UV2/POL 3234 (X-Ray Binaries) The High Speed Photometer was used to make 17 (both photometric and polarimetric) observations of the x-ray binary star HD77581 (aka, Vela XR-1). The visual magnitude of the star is 6.9 . All of the observations were received and displayed in the OSS. All of the observations were reported to look good. 1.6 Completed FOS/BL 4212 (Internal Wave Length Calibrations) The Faint Object Spectrograph was used to make wave length calibration observations to check for unit's stability. The observations were executed as scheduled, and no problems were reported. 1.7 Initiated GHRS/2 4217 (Absorption Cross Sections of Nearby Galaxies) The High Resolution Spectrograph was used to make spectral observations of quasar 1341+258. The metal lines in the quasar's spectra are thought to arise from matter associated with intervening nearby galaxies. The observations executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 9 Successful : 9 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 8 Successful : 8 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete GHRS/2 4217 (Absorption Cross sections of Nearby Galaxies) 3.2 Complete PC 4084 (Globular Cluster Cores, GTO Augmentation) 3.3 Complete FOC/48 4740 (Intensifier HV Test) 3.4 Complete PC 4140 (Rapid Internal Monitor) 3.5 Complete Two Sets of HSP/UV2/POL 3234 (X-Ray Binaries) 3.6 Complete HSP/UV2 3255 (Optical Variability Associated with Black Holes) 3.7 Complete WFC 1306 (GTO Low Latitude Parallel Program) 3.8 Complete FOS/RD 4212 (Internal Wave Length Calibration) 3.9 Complete Two Sets of FOC/96 4107 (Search for QSOs Suitable for Subsequent Observations) 3.10 Complete FOC/96 3801 (Search to QSOs Suitable for Observation) 3.11 Complete FOC/96 4075 (Magellanic Cloud Planetary Nebulae) 3.12 Complete FOS/BL 4110 (The Physics of Massive O Stars) 3.13 Complete WFC 1297 (GTO High Latitude Parallel Program) 3.14 Complete WFC 4105 (HST Medium Deep Survey: High Latitude) 3.15 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.14 Complete WFC-PC 4109 (Rapid Internal Monitor) 3.15 Complete FGS-PC 4765 (Overlapping Plate Field Distortion) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time -END-